### v2

@@ 1 Introduction
Variational image reconstruction has proven to be a successful approach for solving inverse problems in imaging such as image reconstruction. The reconstruction can then be obtained by solving the following minimization problem:
@@
$\min_{\theta}\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{x}_{i}(\theta))$ (2)
$\hat{x}_{i}(\theta):=\arg\min_{x}\Phi_{i}(x;\theta),\ i=1,\dots,n,$
where the reconstruction $\hat{x}_{i}(\theta)$ is the solution to the lower-level optimization problem, and $\ell$ is a function that quantifies the quality of the reconstruction, such as the mean squared error $\|\hat{x}_{i}(\theta)-x^{*}_{i}\|^{2}$. Examples of learnable parameters $\theta$ include feature-extracting filters, wavelet- or shearlet-based systems for regularizers.
@@ 2.1 Problem Setting
For the lower-level problem in (2), we consider the general optimization problem:
$\min_{x\in\mathcal{X}}\{\Phi(x;K):=f(Kx)+g(x)\},$ (P)
where $f$ and $g$ are proper, convex, lower semicontinuous functions. The corresponding dual problem is given by:
@@ 2.2 Piggyback Algorithm
It was shown in [28, Theorem 2.2] that under 1, 2, and 3, and with a suitable choice of parameters for PDHG, i.e. step-sizes $\tau,\sigma$ and extrapolation factor $\theta$, this alternative piggyback algorithm converges linearly to $(\hat{x}(K),\hat{y}(K),\hat{X}(K),\hat{Y}(K))$, and thus to the hypergradient $\nabla\mathcal{L}(K)$.
@@ 3.1 Inexact Piggyback Algorithm and A-Posteriori Error Analysis
$\leq(L_{\nabla g})^{2}\|\nabla^{2}g^{*}(y_{1})-\nabla^{2}g^{*}(y_{2})\|,$
where the last inequality is due to [43, Theorem 1] and the $L_{\nabla g}$ smoothness of $g$. Since $g^{*}$ is $C^{2,1}$ and $g$ is $L_{\nabla g}$ smooth, we have:
@@
Taking similar steps on (AS) and defining
@@
where $(\hat{X},\hat{Y})$ is the saddle-point of (AS). Now, from the RHS of (19), we can write:
@@
where the inequality is due to the $L_{1}$-smoothness of $\ell_{1}$. Note that we have:
@@
$=L_{\nabla^{2}g^{*}}(L_{\nabla g})^{3}\epsilon^{x}+L_{\nabla^{2}f}L_{\nabla f^{*}}\|K\|^{2}\epsilon^{y}.$
We note that:
@@
where the last inequality again utilizes [43, Theorem 1]. Moreover, we have:
@@ 3.2 Method of Adaptive Inexact Descent (MAID)
$\psi(\alpha_{k})\coloneqq\ell(u_{t+1})+\|\nabla\ell(u_{t+1})\|\bar{\epsilon}\\ +\frac{L_{1}+L_{2}}{2}\bar{\epsilon}^{2}-\ell(u_{t})+\|\nabla\ell(u_{t})\|\bar{\epsilon}+\lambda\alpha_{t}\|z_{t}\|^{2}.$ (27)
@@ 4.1 Learning the Discretization of Total Variation
For a given set of corresponding ground truth images and corrupted observations $(x_{i}^{*},u_{i})_{i}$, the standard bilevel setting in (4) can be employed, which given the saddle-point structure of the underlying lower-level problem can be approached using Algorithm 3. The experiments were conducted for the inverse problem of image denoising for a randomly sampled train subset of the BSDS train data set [51] with 32 images of size $128\times 128$. To generate the corrupted observations for the image denoising task, zero-mean additive Gaussian noise $\sim\mathcal{N}(0,\sigma^{2})$ with $\sigma=25.5$ was used.
@@
The linear operator $K$ was heuristically chosen to consist of $n=8$ filters with filter kernels $5\times 5$. The parameters in Algorithm 3 were set to $\underline{\nu}=0.5$, $\bar{\nu}=1.05$, $\underline{\rho}=0.5$, $\bar{\rho}=\tfrac{10}{9}$ following [39]. For comparison between the adaptive and non-adaptive settings, where the latter corresponding to the native inexact version of piggyback, the initial values for $\epsilon_{0}$, $\delta_{0}$ and $\alpha_{0}$ were kept the same across both cases. However, in the non-adaptive version, these parameters were fixed throughout the algorithm. Figure 1(a) shows the learned filters with the adaptive Algorithm 3 and without for $\epsilon_{0}=10^{-3}$, $\delta_{0}=10^{-3}$ and $\alpha_{0}=10^{-2}$, showing that both settings seem to capture very related orientations. While clearly the quantitatively learned filters are very similar, it is of higher interest to investigate the resulting implications on the used computational budget in terms of required lower-level iterations. Therefore, Figure 1(b) shows the resulting upper-level loss depending on the required total computational budget in all standard and adjoint primal-dual algorithms for different experimental settings. The benefit of adaptively choosing the step-size can clearly be seen as the resulting upper-level loss decreases smoothly with increasing computational budget in the lower-level solver. On the other hand, the dashed curves show scenarios with the same initial accuracies and learning rate $\alpha$ as in the adaptive setting, but these parameters are now kept constant throughout the learning algorithm. Depending on the initial learning rate, the upper-level loss even starts oscillating after an initial phase of decrease.
@@ 4.2 Learning Convex Regularizers
Figure 4 shows the qualitative results of reconstructing the test image in the sparse-view measurement scenario for all the above described methods as well as the corresponding ground truth image. Visually, the reconstruction results of the 2-layer ICNN trained using our proposed bilevel learning approach, as well as those from the learned primal-dual (LPD) method, exhibit high-quality reconstructions. They avoid both oversmoothed artifacts and overly fine-grained structures, which can often occur with other reconstruction methods. To support these results, we also show quantitative results in Table 1, which contains peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) for all methods on the employed test phantom. Note that this underlines the qualitative impressions from Figure 4. While the best scores are obtained using LPD, also the ICNN regularizer trained in our bilevel settings shows very competitive results. To some extent this can also be explained the number of learnable parameters which is in a ten-fold increase for LPD. On the other hand, the Fields of Experts regularizer that was trained using Algorithm 3 shows remarkable reconstruction results given its reduced complexity and limited number of learnable parameters. Even more interestingly, using the exact same architecture for the 2-layer ICNNs, our bilevel learning framework manages to outperform ACR by a large margin. Since they have identical learnable parameters, this shows very well the efficiency of the employed bilevel learning framework using primal-dual style differentiation. Overall the experiments show competitive results of FoE and ICNN-based regularizers trained using Algorithm 3. Future work could investigate the performance of the proposed method in more challenging scenarios, such as limited-angle measurement settings. Furthermore, the entire framework can be directly extended to other linear inverse problems with data-driven and potentially non-smooth regularizers.
Table 1: PSNR and SSIM values of test reconstructions shown in Figure 4 for different methods for sparse-view CT reconstruction. Best score for each metric in bold, second-best underlined.

### v3

@@ 1 Introduction
Variational image reconstruction has proven to be a successful approach for solving inverse problems in imaging such as image reconstruction. The reconstruction can be obtained by solving the following minimization problem:
@@
$\min_{\theta}$$\left\{\mathcal{L}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{x}_{i}(\theta))\right\}$ (2)
s.t. $\hat{x}_{i}(\theta)=\arg\min_{x}\Phi_{i}(x;\theta),$
where $\Phi_{i}(x;\theta)\coloneqq D(Ax,u_{i})+R_{\theta}(x)$ denotes the lower-level objective function for each data sample $i=1,\dots,n$, the reconstruction $\hat{x}_{i}(\theta)$ is the solution to the lower-level optimization problem with respect to measurement $u_{i}$, and $\ell$ is a function that quantifies the quality of the reconstruction, such as the mean squared error $\|\hat{x}_{i}(\theta)-x^{*}_{i}\|^{2}$.
We focus on the setting where the learnable parameter $\theta$ is a linear operator $K$ and the lower-level problem in (2) can be formulated as the following general optimization problem:
$\min_{x\in\mathcal{X}}\{\Phi(x;K):=f(Kx)+g(x)\},$ (P)
where $f$ and $g$ are proper, convex, lower semicontinuous functions. For example, this corresponds to $R_{\theta}(x)=\psi(Wx)$ in the Fields of Experts model, with the filters $W$ and the activation $\psi$ acting as the learnable operator $K$ and the function $f$ respectively, while the data fidelity acts as $g$. In the setting of training an input-convex neural network, (P) corresponds to a reformulation of the lower-level problem in (2), which will be presented in later section.
@@ 2.1 Problem Setting
The corresponding dual problem of (P) is given by:
@@ 2.2 Piggyback Algorithm
It was shown in [28, Theorem 2.2] that under 1, 2, and 3, and following the parameter selection condition on the step-sizes $\tau,\sigma$, and extrapolation factor $\theta$ of PDHG [42, Algorithm 3], this alternative piggyback algorithm converges linearly to $(\hat{x}(K),\hat{y}(K),\hat{X}(K),\hat{Y}(K))$, and thus to the hypergradient $\nabla\mathcal{L}(K)$.
@@ 3.1 Inexact Piggyback Algorithm and A-Posteriori Error Analysis
$\leq(L_{\nabla g})^{2}\|\nabla^{2}g^{*}(y_{1})-\nabla^{2}g^{*}(y_{2})\|.$
For the first inequality, we utilized the fact that $\nabla^{2}g(x_{i})^{-1}=\nabla^{2}g^{*}(y_{i})$ for $i=1,2$, and the last inequality is due to [43, Theorem 1] and the $L_{\nabla g}$ smoothness of $g$. Since $g^{*}$ is $C^{2,1}$ and $g$ is $L_{\nabla g}$ smooth, we have:
@@
Taking similar steps as Lemma 3 on (AS) and defining
@@
where $(\hat{X},\hat{Y})$ is the saddle-point of (AS). Now, from the RHS of (19), by adding and subtracting $\tilde{B}\tilde{X}+\nabla_{1}(\tilde{x})$, we have:
@@
where the inequality is due to the $L_{1}$-smoothness of $\ell_{1}$ and the triangle inequality. Note that we have:
@@
$=L_{\nabla^{2}g^{*}}(L_{\nabla g})^{3}\epsilon^{x}+L_{\nabla^{2}f}L_{\nabla f^{*}}\|K\|^{2}\epsilon^{y},$
where the first inequality is due to Lemma 4 and the Lipschitz condition on $\nabla^{2}f$, and the last inequality is due to 1 and Remark 1. We note that:
@@
where the last inequality again utilizes [43, Theorem 1]. Moreover, since $\nabla^{2}f^{*}(y)^{-1}=\nabla^{2}f(\nabla f^{*}(y))$ for all $y$, we have:
@@
Remark 3.
For a general $\ell$, $\nabla\ell_{1}$ and $\nabla\ell_{2}$ will be replaced by $\nabla_{x}\ell$ and $\nabla_{y}\ell$ in Lemma 3 and Theorem 2 respectively. In the proof of Theorem 2, error bounds on the gradients of upper-level loss such as $\|\nabla_{1}(\hat{x})-\nabla_{1}(\tilde{x})\|$ can be bounded by $L(\epsilon^{x}+\epsilon^{y})$ instead.
@@ 3.2 Method of Adaptive Inexact Descent (MAID)
$\psi(\alpha_{t})\coloneqq\ell(u_{t+1})+\|\nabla\ell(u_{t+1})\|\bar{\epsilon}\\ +\frac{L_{1}+L_{2}}{2}\bar{\epsilon}^{2}-\ell(u_{t})+\|\nabla\ell(u_{t})\|\bar{\epsilon}+\lambda\alpha_{t}\|z_{t}\|^{2}.$ (27)
@@ 4.1 Learning the Discretization of Total Variation
For a given set of corresponding ground truth images and corrupted observations $(x_{i}^{*},u_{i})_{i}$, the standard bilevel setting in (4) can be employed, which given the saddle-point structure of the underlying lower-level problem can be approached using Algorithm 3. The experiments were conducted for the inverse problem of image denoising for a randomly sampled train subset of the BSDS train data set [51] with 32 images of size $128\times 128$. To generate the corrupted observations for the image denoising task, zero-mean additive Gaussian noise $\sim\mathcal{N}(0,\sigma^{2})$ with $\sigma=25.5$ was used for image intensities scaled between 0 and 255. To evaluate the learned filters, another set of 32 test images of the same size and with the noise level was generated.
@@
The linear operator $K$ was heuristically chosen to consist of $n=8$ filters with filter kernels $5\times 5$. Across all experiments, filter weights were initialized with He-uniform initialization [52] and a constant random seed. The parameters in Algorithm 3 were set to $\underline{\nu}=0.5$, $\bar{\nu}=1.05$, $\underline{\rho}=0.5$, $\bar{\rho}=\tfrac{10}{9}$ following [39]. For comparison between the adaptive and non-adaptive settings, where the latter corresponding to the native inexact version of piggyback, the initial values for $\epsilon_{0}$, $\delta_{0}$ and $\alpha_{0}$ were kept the same across both cases. However, in the non-adaptive version, these parameters were fixed throughout the algorithm. Figure 1(a) shows the learned filters with the adaptive Algorithm 3 and without for $\epsilon_{0}=10^{-3}$, $\delta_{0}=10^{-3}$ and $\alpha_{0}=10^{-2}$, showing that both settings seem to capture very related orientations. While clearly the quantitatively learned filters are very similar, it is of higher interest to investigate the resulting implications on the used computational budget in terms of required lower-level iterations. Therefore, Figure 1(b) shows the resulting upper-level loss depending on the required total computational budget in all standard and adjoint primal-dual algorithms for different experimental settings. The benefit of adaptively choosing the step-size can clearly be seen as the resulting upper-level loss decreases smoothly with increasing computational budget in the lower-level solver. On the other hand, the dashed curves show scenarios with the same initial accuracies and learning rate $\alpha$ as in the adaptive setting, but these parameters are now kept constant throughout the learning algorithm. Depending on the initial learning rate, the upper-level loss even starts oscillating after an initial phase of decrease.
@@ 4.2 Learning Convex Regularizers
Figure 4 shows the qualitative results of reconstructing the test image in the sparse-view measurement scenario for all the above described methods as well as the corresponding ground truth image. Visually, the reconstruction results of the 2-layer ICNN trained using our proposed bilevel learning approach, as well as those from the learned primal-dual (LPD) method, exhibit high-quality reconstructions. They avoid both oversmoothed artifacts and overly fine-grained structures, which can often occur with other reconstruction methods. To support these results, we also show quantitative results for the test image in Table 1, which contains peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) for all methods on the employed test phantom. Note that this underlines the qualitative impressions from Figure 4. While the best scores are obtained using LPD, also the ICNN regularizer trained in our bilevel settings shows very competitive results. To some extent this can also be explained the number of learnable parameters which is in a ten-fold increase for LPD. On the other hand, the Fields of Experts regularizer that was trained using Algorithm 3 shows remarkable reconstruction results given its reduced complexity and limited number of learnable parameters. Even more interestingly, using the exact same architecture for the 2-layer ICNNs, our bilevel learning framework manages to outperform ACR by a large margin. Since they have identical learnable parameters, this shows very well the efficiency of the employed bilevel learning framework using primal-dual style differentiation. Overall the experiments show competitive results of FoE and ICNN-based regularizers trained using Algorithm 3. Future work could investigate the performance of the proposed method in more challenging scenarios, such as limited-angle measurement settings. Furthermore, the entire framework can be directly extended to other linear inverse problems with data-driven and potentially non-smooth regularizers.
Table 1: PSNR and SSIM values of test reconstructions shown in Figure 4 for different methods for sparse-view CT reconstruction on the test image. Best score for each metric in bold, second-best underlined.